If you need 25,000 A of current to carry out electrolysis for 15 minutes, what voltage would you need if you are to use 1.7 kWh for the process?


1
Expert's answer
2021-06-04T09:00:01-0400

Solution


E=QV

Q=I"\times" t

t=15"\times" 60=900 sec

2500"\times" 900

Q=22.5"\times10"6

E=P"\times" t

converting 1.7kwh to joules

E=1.7"\times3.6\times10^6\times900"

E=QV

therefore V="\frac{E}{V}"


V="\frac{1.7\times{3.6}\times{10^6}\times{900}}{22.5\times{10^6}}"


v=244.5